A time keeper expressed 1/3 of an hour in minute :instead of 1/4 of it. Calculate the error in percent
Answer by stanbon(75887)

A time keeper expressed 1/3 of an hour in minute :instead of 1/4 of it. Calculate the error in percent
-----
Quantitative error:: (1/3)-(1/4) = (4-3)/12 = 1/12 minute
-----
Percentage error:: (1/12)/(1/4) = 1/3 = 33 1/3%
-----------
Cheers,
Stan H.
-----------------